OpenStudy (anonymous):

this one ? @tHe_FiZiCx99 7x+3 = x+3

OpenStudy (anonymous):

trying to get into slope intercept from

OpenStudy (jadeishere):

7x + 3 = x + 3 7x = x + 0 hm...

OpenStudy (anonymous):

lol, that is the question.

OpenStudy (anonymous):

these are the choices

OpenStudy (jadeishere):

/).^

OpenStudy (the_fizicx99):

Plug in x. 7(-1) + 3 = -1 + 3 -7 + 3 = -4 -4 = 2 <- not true. Look for the one that will give you -4 = -4

OpenStudy (anonymous):

is it the third one ?

OpenStudy (anonymous):

-4 = -4

OpenStudy (the_fizicx99):

ye

OpenStudy (anonymous):

thanks

OpenStudy (the_fizicx99):

yw